​A 6.96-nC charge is located 1.65 m from a 4.06-nC point charge.


A 6.96-nC charge is located 1.65 m from a 4.06-nC point charge. 

(a) Find the magnitude of the electrostatic force that one charge exerts on the other. 

(b) Is the force attractive or repulsive? 
Answer #1

here,

charge 1 , q1 = 6.96 nC = 6.96 * 10^-9 C

charge 2 , q2 = 4.06 nC = 4.06 * 10^-9 C

sepration between charges , r = 1.65 m

a)

the magnitude of the electrostatic force that one charge exerts on the other , F = K * q1 * q2 /r^2

F = 9 * 10^9 * 6.96 * 10^-9 * 4.06 * 10^-9 /1.65^2

F = 9.34 * 10^-8 N

b)

as both charges have same sign

the force is repulsive